# Pinpointing Important Genetic Variants via A Feature-level Filter and Group Knockoffs

### v2

@@ 4 AD Analysis of EADB–UKBB Dataset
Given the promising performance of the proposed filter in simulated experiments, we apply our method to the EADB-UKBB dataset where model-X knockoff filter exhibits power loss and the group knockoff filter returns not informative enough results in Section 1.2.
To infer which variants are associated with AD, we first compute correlation $\text{cor}(G_{i},G_{j})$ between any pair of variants $G_{i}$ and $G_{j}$ and construct variant groups $B_{1},\ldots,B_{K}$ by applying the hierarchical clustering (average linkage with cutoff value $0.5$) on the distance matrix $(1-|\text{cor}(G_{i},G_{j})|)_{p\times p}$. By doing so, we obtain the same $321,535$ groups $B_{1},\ldots,B_{k}$ as the group knockoff filter in Section 1.2 where the average absolute correlation of variants from different groups are lower than 0.5. Based on the correlation matrix $(\text{cor}(G_{i},G_{j}))_{p\times p}$ and variant groups $B_{1},\ldots,B_{K}$, we generate group knockoffs $\widetilde{\mathbb{G}}$ under the maximum entropy (ME) construction of Chu et al., (2023). Given the response of interest $\textbf{y}=(y_{1},\ldots,y_{n})^{\mathsf{T}}$ where $y_{i}=1$ if the $i$-th observation corresponds to a clinically diagnosed AD case and $y_{i}=0$ otherwise, we use absolute values of lasso estimators of the logistic regression model,
@@
as feature importance scores. In other words, we compute $T_{j}=\left|\widehat{\beta}_{j}\right|$ and $\widetilde{T}_{j}=\left|\widehat{\widetilde{\beta}}_{j}\right|$, which are also used in other knockoff-based methods and provide more informative rejection sets without large power loss given the large sample size. We then calculate feature statistics $W_{j}=\text{sign}\left(T_{j}-\widetilde{T}_{j}\right)\cdot\max\left(T_{j},\widetilde{T}_{j}\right)$ and implement the FVG filter with target FDR level $\alpha=0.10$. Variants in the rejection set $\mathcal{R}^{(\text{fg})}$ under target FDR level $\alpha=0.10$ are presented in Figure 3 with their post-selection importance scores $W_{j}\cdot\text{I}(W_{j}>0)$ and names of the closest genes.
@@
Details of variants identified by the FVG filter under the target FDR level $0.10$ are provided in Table 3. In total, we identify 205 variants that contribute to AD variation substantially, which comes from 154 groups in 84 loci. Similar to the literature, multiple variants are identified in the APOE/APOC region with the strongest association to AD (chromosome 19, positions 44909011 $\sim$ 45912650). In addition, we also manage to identify variants rs744373 and rs2118506 close to gene “BIN1”, variant rs1532278 close to gene “CLU”, variant rs10792832 close to gene “PICALM”, variants rs4277405, rs4309 and rs3730025 close to gene “ACE”, all of which are also reported in He et al., (2022) and Bellenguez et al., (2022). Compared with the model-X knockoff filter, the FVG filter only misses 2 loci identified by the model-X knockoff filter but discover 37 more loci, which suggests that our method would not suffer from the same power loss issue of the model-X knockoff filter.
@@
Table 3: Details of variants identified by the FVG filter under the EADB–UKBB dataset and target FDR level $\alpha=0.10$.
Identified Position Closest Identified Position Closest Identified Position Closest Identified Position Closest
Variant Gene Variant Gene Variant Gene Variant Gene
Chromosome 1 rs3997700 41251889 TREML5P rs9516168 93085490 LINC00363 rs3852860 44879709 NECTIN2
rs11579514 161173522 B4GALT3 rs9296558 47484147 CD2AP Chromosome 14 rs365653 44858389 NECTIN2
rs782791 203029691 PPFIA4 rs10948363 47520026 CD2AP rs74093831 105733666 ELK2AP rs73052307 44881148 NECTIN2
rs4562624 207512620 CR1 rs12525381 696327 PTPRK rs873533 106667442 LINC00221 rs11668327 44895376 TOMM40
rs3818361 207611623 CR1 rs117591349 78478451 MIR10524 rs17125924 52924962 FERMT2 rs449647 44905307 APOE
rs6701713 207612944 CR1 rs313161 85332752 LINC02535 rs4904929 92470949 SLC24A4 rs10420036 44882772 NECTIN2
rs4844610 207629207 CR1 Chromosome 7 Chromosome 15 rs283813 44885917 NECTIN2
Chromosome 2 rs221834 100745552 ZAN rs59685680 50709337 SPPL2A rs7254892 44886339 NECTIN2
rs116038905 105805908 NCK2 rs144867634 111940111 DOCK4 rs28470808 50735410 SPPL2A rs7412 44908822 APOE
rs74851408 117201103 DDX18 rs10952168 1554792 TMEM184A rs2043085 58388755 LIPC rs445925 44912383 APOC1
rs6733839 127135234 BIN1 rs56402156 143406388 TCAF2 rs593742 58753575 ADAM10 rs390082 44913574 APOC1
rs730482 127136908 BIN1 rs12703526 143410495 TCAF2 rs117618017 63277703 APH1B rs190712692 44921921 APOC1
rs744373 127137039 BIN1 rs9640386 143435090 TCAF2 rs3816385 64126456 SNX1 rs41290108 44874585 NECTIN2
rs745717 127136321 BIN1 rs2189965 28132395 JAZF1 rs56117790 79075361 RASGRF1 rs73033507 44928146 APOC1P1
rs2118506 127139927 BIN1 rs2122623 37842715 NME8 Chromosome 16 rs79701229 44881674 NECTIN2
rs10496731 134840058 ACMSD rs7809394 55077689 EGFR rs12325539 30022312 C16orf92 rs157580 44892009 TOMM40
rs4954187 134842848 ACMSD rs12669393 7815748 UMAD1 rs11865499 31120929 KAT8 rs2075649 44892073 TOMM40
rs3819121 134865290 ACMSD rs4725076 8204779 ICA1 rs8058370 81738205 CMIP rs8106922 44898409 TOMM40
rs6725887 202881162 WDR12 rs1476679 100406823 ZCWPW1 rs12446759 81739398 CMIP rs405697 44901434 TOMM40
rs10933431 233117202 INPP5D rs41280968 100105324 AP4M1 rs11548656 81883307 PLCG2 rs7259620 44904531 TOMM40
rs7421448 233117495 INPP5D rs13228694 100342684 STAG3L5P-PVRIG2P-PILRB rs1071644 81937798 PLCG2 rs440446 44905910 APOE
rs35349669 233159830 INPP5D Chromosome 8 rs56407236 90103687 FAM157C rs6857 44888997 NECTIN2
rs3771371 71400409 ZNF638 rs2436889 102564430 ODF1 Chromosome 17 rs2075650 44892362 TOMM40
rs113068192 76786226 LRRTM4 rs4731 11808828 FDFT1 rs79759095 44451394 GPATCH8 rs157581 44892457 TOMM40
Chromosome 3 rs79832570 144042819 SPATC1 rs5910 44372421 ITGA2B rs34404554 44892652 TOMM40
rs7634492 136105288 PPP2R3A rs61732533 144053248 OPLAH rs199439 46716137 NSF rs75627662 44910319 APOE
rs16824536 155069722 MME rs34674752 144099319 SHARPIN rs199451 46724418 NSF rs438811 44913484 APOC1
rs61762319 155084189 MME rs73223431 27362470 PTK2B rs2671659 49397167 PHB rs283814 44885967 NECTIN2
rs4974180 56200107 ERC2 rs1532278 27608798 CLU rs616338 49219935 ABI3 rs439401 44911194 APOE
Chromosome 4 rs9297949 94957217 NDUFAF6 rs2072153 49312652 ZNF652 rs59325138 44913034 APOC1
rs6448453 11024404 CLNK Chromosome 9 rs2526378 58326988 TSPOAP1-AS1 rs584007 44913221 APOC1
rs13116567 12239152 LINC02270 rs1131773 93077974 SUSD3 rs2632516 58331728 TSPOAP1-AS1 rs157595 44922203 APOC1
rs502746 14019593 LINC01182 Chromosome 10 rs9896864 63458947 CYB561 rs4803770 44924096 APOC1P1
rs116791081 143428212 GAB1 rs7920721 11678309 ECHDC3 rs4277405 63471557 ACE rs71352239 44926286 APOC1P1
Chromosome 5 rs7068614 11679478 ECHDC3 rs4309 63482562 ACE rs118060185 44928417 APOC1P1
rs25992 14707491 ANKH rs10764258 20634336 MIR4675 rs3730025 63480412 ACE rs144311893 44920687 APOC1
rs871269 151052827 TNIP1 rs117980033 29966853 JCAD rs9901675 7581494 CD68 rs114533385 44933496 APOC1P1
rs335473 177559423 FAM193B rs1227759 69831131 COL13A1 Chromosome 19 rs79429216 44942260 APOC4-APOC2
rs4700716 180138792 RASGEF1C rs6586028 80494228 TSPAN14 rs445752 43956053 ZNF221 rs12721109 44943964 APOC4-APOC2
rs10068419 86923485 LINC02059 Chromosome 11 rs2571174 44430314 ZNF229 rs16979595 44974124 CLPTM1
Chromosome 6 rs74685827 121482368 SORL1 rs2293164 44475997 ZNF180 rs10411314 45224801 EXOC3L2
rs976271 114361563 HDAC2-AS2 rs2298813 121522975 SORL1 rs1625502 44540845 CEACAM22P rs75987583 45745299 MEIOSIN
rs2130357 27919052 OR2B2 rs3781837 121578263 SORL1 rs10405071 44732632 BCL3 rs875121 49948251 SIGLEC11
rs13408 28244970 ZKSCAN4 rs3740890 130385218 ZBTB44-DT rs2965169 44747899 BCL3 rs1710398 51223655 CD33
rs3130931 31167111 POU5F1 rs17439299 15599040 LINC02751 rs116480598 44753996 BCL3 rs3865444 51224706 CD33
rs3132451 31614248 AIF1 rs1685404 47222114 DDB2 rs28399654 44813331 BCAM rs12974390 54264111 LILRB5
rs532965 32610196 HLA-DRB1 rs2293579 47419207 PSMC3 rs10419669 44794812 CBLC rs10404195 5908958 VMAC
rs9469112 32447376 HLA-DRA rs1582763 60254475 MS4A4E rs10405693 44823407 BCAM rs4147918 1058177 RDH13
rs9271162 32609938 HLA-DRB1 rs11824773 60309467 MS4A4A rs73572039 44839373 NECTIN2 19:1051137_CTG_C 1051138 RDH13
rs9273349 32658092 HLA-DQB1 rs17854357 65834089 SNX32 rs77301115 44893716 TOMM40 rs3752246 1056493 RDH13
rs9275184 32686937 HLA-DQB1 rs536841 86076782 PICALM rs112849259 44894050 TOMM40 rs35917007 1849148 REXO1
rs9275614 32716480 HLA-DQA2 rs10792832 86156833 PICALM rs79398853 44895528 TOMM40 Chromosome 20
rs114812713 41066261 OARD1 rs3844143 86139201 PICALM rs75687619 44896087 TOMM40 rs6064392 56409712 CASS4
rs7748513 41160234 TREM2 Chromosome 12 rs76366838 44896639 TOMM40 rs6069749 56446519 CASS4
rs2093395 41187288 TREML2 rs17860282 51346566 CELA1 rs115881343 44899959 TOMM40 Chromosome 21
rs143332484 41161469 TREM2 Chromosome 13 rs3865427 44877704 NECTIN2 rs62222965 26225443 APP
rs9394778 41247320 TREML5P
Figure 4: Comparison of our FVG filter and the group knockoff filter in analyzing the EADB–UKBB dataset: (a) Venn diagram of variant groups identified by different filters; (b) bar plot of size of catching sets identified by different filters (columns); and (c) histogram of purity of catching sets identified by different filters.
Compared to the group knockoff filter which identifies 152 AD-associated variant groups from 84 loci under the same target FDR level $0.10$, our FVG filter does not suffer power loss. As shown in Figure 4 (a), among all variant groups identified by either the group knockoff filter or our FVG filter [^4] Here, we say a variant group is identified by our FVG filter if it contains at least one variant identified by our FVG filter ., 143 groups are identified by both methods while 11 groups and 9 groups are only identified by our filter and the group knockoff filter respectively. This suggests the consistency of inference results between our FVG filter and the group knockoff filter. However, our filter manages to detect signals at variant rs1532278 close to gene “CLU”, while none of variant groups in the “CLU” locus are identified by the group knockoff filter. Without significant power loss, our method is shown to succeed in pinpointing AD-associated variants with higher precision. As shown in Figure 4 (b)-(c), all catching sets obtained by our method are of size no larger than $6$ while most of catching sets (129/154) are of size 1 and purity greater than $90\%$. In other words, most of catching sets only contain the important variants or strongly correlated proxy variants. In contrast, 37 catching sets obtained by the group knockoff filter (Dai and Barber,, 2016) contain more than $8$ variants and are far from being informative with strong impurity. Specifically, for the five largest catching sets obtained by the group knockoff filter shown in Table 4, our FVG filter manages to distinguish one or two variants from more than 100 proxy variants in each. This suggests the ability of the proposed method in pinpointing AD-associated variants. Comprehensive comparisons of our method with all existing methods are presented in Table 5.
@@
Identified Chromosome Position Catching sets obtained by Closest
@@
rs199439 17 46716137 17:43056905:45876021:36 (314) NSF
rs199451 17 46724418
Table 5: Summary of results by applying different methods on the EADB-UKBB dataset.
@@
Marginal association test 51 18.078 18.078 0.477
Model-X knockoff filter 49 2.816 1.000 1.000
Group knockoff filter 84 23.714 13.105 0.610
FVG filter 84 2.440 1.257 0.958

### v3

@@ 4 AD Analysis of EADB–UKBB Dataset
Given the promising performance of the proposed filter in simulated experiments, we apply it to the EADB-UKBB dataset where model-X knockoff filter and KeLP exhibits power loss and the group knockoff filter and KnockoffZoom returns not informative enough results in Section 1.2.
To infer which variants are associated with AD, we first compute correlation $\text{cor}(G_{i},G_{j})$ between any pair of variants $G_{i}$ and $G_{j}$ using the UK Biobank directly genotyped data as the reference panel and construct variant groups $B_{1},\ldots,B_{K}$ by applying the hierarchical clustering (average linkage with cutoff value $0.5$) on the distance matrix $(1-|\text{cor}(G_{i},G_{j})|)_{p\times p}$. By doing so, we obtain the same $321,569$ groups $B_{1},\ldots,B_{k}$ as the group knockoff filter in Section 1.2 where the average absolute correlation of variants from different groups are lower than 0.5. Based on the correlation matrix $(\text{cor}(G_{i},G_{j}))_{p\times p}$ and variant groups $B_{1},\ldots,B_{K}$, we generate second-order group knockoffs $\widetilde{\mathbb{G}}$ under the maximum entropy (ME) construction using Algorithm 2 of Chu2024 by assuming that the distribution of X can be well approximated by a multivariate Gaussian distribution. Given the response of interest $\textbf{y}=(y_{1},\ldots,y_{n})^{\mathsf{T}}$ where $y_{i}=1$ if the $i$-th observation corresponds to a clinically diagnosed AD case and $y_{i}=0$ otherwise, we use absolute values of lasso estimators of the logistic regression model,
@@
as feature importance scores. In other words, we compute $T_{j}=\left|\widehat{\beta}_{j}\right|$ and $\widetilde{T}_{j}=\left|\widehat{\widetilde{\beta}}_{j}\right|$, which are also used in other knockoff-based methods and provide more informative rejection sets without large power loss given the large sample size. We then calculate feature statistics $W_{j}=\text{sign}\left(T_{j}-\widetilde{T}_{j}\right)\cdot\max\left(T_{j},\widetilde{T}_{j}\right)$ and implement the FVG filter with target FDR level $\alpha=0.10$.
@@
Table 3: Details of variants identified by the FVG filter under the EADB–UKBB dataset and target FDR level $\alpha=0.10$. Here, the column “cS2Gene” denotes the gene that each genetic variant functionally annotates, and the column “Annotation” depicts the type of annotation.
Identified Position cS2Gene Annotation Identified Position cS2Gene Annotation Identified Position cS2Gene Annotation
Variant Variant Variant
Chromosome 1 rs3740890 130385218 rs419010 44865063 PVRL2 GTeX
rs2305463 20853688 EIF4G3 Exon Chromosome 12 rs365653 44858389
rs782791 203029691 PPFIA4 GTeX,ABC rs7183 7070715 C1S Exon rs17561351 44869072 TOMM40 eQTLGen,ABC
rs4844610 207629207 CR1 GTeX rs17860282 51346566 CELA1 Promoter rs73052307 44881148 APOE EpiMap
Chromosome 2 Chromosome 13 rs11668327 44895376
rs35933838 9544047 CPSF3 eQTLGen rs9516168 93085490 rs449647 44905307 APOE Promoter,ABC
rs6726709 37270395 Chromosome 14 rs41290122 44880326
rs77059113 43445369 rs17125924 52924962 rs10420036 44882772 APOE EpiMap
rs113068192 76786226 rs12590654 92472511 rs283808 44883777 PVRL2 GTeX
rs116038905 105805908 rs4904929 92470949 rs283813 44885917 PVRL2 Exon
rs74851408 117201103 rs74093831 105733666 CRIP1 EpiMap rs7254892 44886339 APOE EpiMap=0.500
rs6733839 127135234 ERCC3 Cicero rs873533 106667442 rs7412 44908822 APOE Exon
rs744373 127137039 BIN1 GTeX Chromosome 15 rs445925 44912383 APOE GTeX
rs2118506 127139927 BIN1 EpiMap rs12595082 50715532 rs390082 44913574 APOC1 Promoter,EpiMap
rs4954187 134842848 rs2043085 58388755 LIPC GTeX rs190712692 44921921
rs6725887 202881162 rs593742 58753575 rs141622900 44923535
rs10933431 233117202 NGEF GTeX rs16946801 63312881 CA12 GTeX rs41290100 44867684
rs7421448 233117495 INPP5D EpiMap rs117618017 63277703 APH1B Exon,GTeX,eQTLGen,ABC rs187183066 44872328 CEACAM22P GTeX
Chromosome 3 rs56117790 79075361 rs41290108 44874585 PVRL2 Promoter
rs4974180 56200107 Chromosome 16 rs73033507 44928146 APOC4 EpiMap
rs7634492 136105288 SLC35G2 GTeX rs12325539 30022312 DOC2A Promoter rs79149284 45010596 DMWD GTeX
rs16824536 155069722 rs11865499 31120929 KAT8 GTeX,eQTLGen rs77196615 44877078 ZNF296 EpiMap
rs61762319 155084189 MME Exon rs78924645 31143037 PRSS36 Promoter,GTeX rs79701229 44881674 PVRL2 Promoter
Chromosome 4 rs8058370 81738205 PLCG2 Promoter rs3745150 44882502 TOMM40 GTeX
rs6448453 11024404 rs12446759 81739398 PLCG2 Promoter,Cicero rs157580 44892009 TOMM40 Promoter
rs13116567 12239152 rs11548656 81883307 PLCG2 Exon rs2075649 44892073 TOMM40 Promoter,GTeX,eQTLGen
rs502746 14019593 rs1071644 81937798 PLCG2 Exon rs8106922 44898409
rs116791081 143428212 rs56407236 90103687 rs405697 44901434 TOMM40 Exon
Chromosome 5 Chromosome 17 rs7259620 44904531 APOC1 GTeX
rs25992 14707491 ANKH Exon,Cicero rs7225151 5233752 SCIMP GTeX,ABC,Cicero rs405509 44905579 APOE Promoter,ABC
rs10068419 86923485 rs9901675 7581494 CD68 Exon rs440446 44905910 APOE Exon,ABC
rs7268 140332965 HBEGF Exon rs79759095 44451394 rs142042446 44883210 APOE EpiMap
rs335473 177559423 RAB24 EpiMap rs199451 46724418 KANSL1-AS1 GTeX rs12972970 44884339 TOMM40 ABC
rs4700716 180138792 RASGEF1C Promoter rs616338 49219935 ABI3 Exon,EpiMap rs283815 44887076
Chromosome 6 rs2632516 58331728 BZRAP1 ABC rs6857 44888997 PVRL2 Exon
rs78799498 17319361 RBM24 ABC rs9896864 63458947 rs71352238 44891079 TOMM40 Promoter,GTeX,ABC
rs2130357 27919052 PGBD1 GTeX rs4277405 63471557 ACE GTeX rs184017 44891712 TOMM40 Promoter,ABC
rs3132451 31614248 AIF1 Promoter,EpiMap rs4309 63482562 ACE Exon,EpiMap rs2075650 44892362
rs532965 32610196 HLA-DRB5 ABC rs3730025 63480412 ACE Exon rs157581 44892457 TOMM40 Exon
rs9469112 32447376 Chromosome 19 rs34095326 44892587 BCAM GTeX
rs9271162 32609938 HLA-DRB1 EpiMap rs4147918 1058177 ABCA7 Exon rs34404554 44892652
rs9273349 32658092 19:1051137_CTG_C 1051138 ABCA7 Exon,GTeX,EpiMap,ABC rs157582 44892962 TOMM40 Promoter
rs9275184 32686937 HLA-DQA2 EpiMap rs3752246 1056493 ABCA7 Exon,GTeX rs10119 44903416 TOMM40 Exon,GTeX
rs3997854 32715138 TAP2 GTeX rs35917007 1849148 REXO1 Promoter,GTeX rs769449 44906745 APOE Promoter,ABC
rs114812713 41066261 OARD1 Exon rs10404195 5908958 VMAC Exon rs75627662 44910319 SNRPD2 GTeX
rs7748513 41160234 TREML2 GTeX rs60184524 43790063 rs7256200 44912678 APOC1 GTeX,EpiMap,ABC
rs2093395 41187288 TREML2 EpiMap rs445752 43956053 rs438811 44913484 APOC1 Promoter,ABC
rs143332484 41161469 TREM2 Exon,EpiMap rs2125579 44288548 ZNF235 Exon,GTeX,eQTLGen rs12721046 44917997
rs9394778 41247320 TREML2 EpiMap rs8107530 44332247 rs12721051 44918903 APOC1 Exon
rs3997700 41251889 rs62116778 44389907 rs111789331 44923868
rs2171089 47548075 CD2AP GTeX rs151111866 44387243 ZNF285 Exon rs73936968 44892559
rs117591349 78478451 rs2571174 44430314 ZNF229 Exon rs439401 44911194 APOE GTeX
rs313161 85332752 rs17800789 44510962 rs59325138 44913034 APOC1P1 GTeX,Cicero
rs976271 114361563 rs7254925 44513811 rs584007 44913221 APOE GTeX
Chromosome 7 rs74602839 44481157 rs157595 44922203
rs10952168 1554792 TMEM184A Promoter,GTeX,ABC rs62116891 44524436 rs1064725 44919304 APOC1 Exon
rs12669393 7815748 rs930461 44539272 CEACAM22P Exon rs118060185 44928417 ZNF155 GTeX
rs3173615 12229791 TMEM106B Exon rs16979269 44552920 ZNF180 GTeX rs144311893 44920687
rs2189965 28132395 rs56757528 44561286 rs4803771 44924391 EML2 GTeX
rs4917014 50266267 IKZF1 EpiMap,Cicero rs10405030 44565601 rs60049679 44926451 APOC1P1 Promoter
rs1476679 100406823 LAMTOR4 GTeX rs2357100 44595918 rs114533385 44933496 ZNF226 GTeX
rs41280968 100105324 AP4M1 Exon,ABC rs7259313 44600383 rs79429216 44942260 APOC4 Exon,GTeX,EpiMap
rs221834 100745552 UFSP1 Cicero rs17658507 44617986 PVR GTeX rs5167 44945208 APOC4 Exon
rs144867634 111940111 DOCK4 Exon rs80122548 44618415 rs3760627 44953923 CLPTM1 Promoter
rs12703526 143410495 EPHA1-AS1 ABC rs62119261 44618732 APOC4 EpiMap rs16979595 44974124
rs7805776 143427203 TAS2R60 eQTLGen rs203717 44636512 IGSF23 Exon rs73560271 45051621 ZNF296 EpiMap
Chromosome 8 rs2965160 44692865 CEACAM19 eQTLGen rs78620885 45087826 PPP1R37 ABC
rs4731 11808828 FDFT1 Exon,ABC rs2965164 44698782 CEACAM16 Promoter rs117106519 45047229 CLASRP ABC
rs73223431 27362470 PTK2B GTeX,eQTLGen rs62120574 44711506 rs3178166 45090912 GEMIN7 Exon,GTeX,eQTLGen
rs9297949 94957217 NDUFAF6 Promoter rs76419583 44714590 CEACAM22P GTeX rs10405086 45123977
rs1693551 100663356 SNX31 Exon,GTeX rs2965109 44722081 BCL3 ABC rs2532037 45194545
rs2436889 102564430 rs1004165 44728939 rs3848526 45195982 SYMPK GTeX
rs79832570 144042819 PARP10 GTeX rs2965169 44747899 BCL3 Exon,ABC rs10411314 45224801 EXOC3L2 Exon
rs61732533 144053248 OPLAH Exon rs116480598 44753996 BCL3 ABC rs17356664 45237513
rs34674752 144099319 SHARPIN Exon,ABC=0.500 rs114294368 44789899 rs77740243 45248200
Chromosome 9 rs75599500 44740439 BCL3 EpiMap rs11559024 45317925 CKM Exon
rs1131773 93077974 SUSD3 Exon,GTeX,EpiMap rs2965174 44741758 CBLC GTeX rs17875640 45323890 CKM Promoter,ABC
rs1883025 104902020 rs10408993 44802496 rs73047784 45458317
Chromosome 10 rs117326714 44783846 rs3916883 45352060 ERCC2 ABC
rs7920721 11678309 USP6NL EpiMap,ABC,Cicero rs28399654 44813331 BCAM Exon rs4802253 45401737 PPP1R13L ABC
rs7068614 11679478 rs117142879 44802327 rs2336219 45409148 PPP1R13L GTeX
rs117980033 29966853 rs28399637 44820881 BCAM Promoter,GTeX,EpiMap,ABC rs7256865 45515149 RTN2 eQTLGen
rs1227759 69831131 COL13A1 GTeX,eQTLGen rs10405693 44823407 BCAM EpiMap rs73942917 45543797 DMWD Cicero
rs6586028 80494228 TSPAN14 GTeX,ABC rs149529419 44868071 rs1264226 45559909 OPA3 GTeX
Chromosome 11 rs140684051 44896199 rs73568973 45549792 OPA3 Exon
rs1685404 47222114 rs73572039 44839373 PVRL2 GTeX rs73570937 45588621 GPR4 EpiMap
rs1582763 60254475 rs76692773 44890954 TOMM40 Promoter,ABC rs75987583 45745299 DMWD GTeX
rs11824773 60309467 MS4A6E Cicero rs61642202 44842026 ZNF224 GTeX rs875121 49948251
rs536841 86076782 rs117310449 44890259 TOMM40 Promoter,ABC rs2075803 51125272 SIGLEC9 Exon,eQTLGen
rs10792832 86156833 PICALM ABC rs77301115 44893716 TOMM40 Promoter,eQTLGen rs1710398 51223655 CD33 GTeX
rs3844143 86139201 EED EpiMap rs79398853 44895528 rs3865444 51224706 CD33 Promoter
rs74685827 121482368 SORL1 ABC,Cicero rs115881343 44899959 TOMM40 eQTLGen rs731170 54664811 LILRB4 Exon
rs2298813 121522975 SORL1 Exon rs41289512 44848259 CLASRP Cicero Chromosome 20
rs3781837 121578263 rs12691088 44915229 APOC1 Exon,ABC rs6064392 56409712 CASS4 GTeX,eQTLGen,EpiMap,ABC
rs12272618 121589615 SORL1 Promoter rs3865427 44877704 PVRL2 GTeX rs6069749 56446519 CASS4 GTeX
Variants in the rejection set $\mathcal{R}^{(\text{fg})}$ under target FDR level $\alpha=0.10$ are presented in Figure 5 with their positive part of importance scores $W_{j}\cdot\text{I}(W_{j}>0)$’s and names of the closest genes. Details of these variants are provided in Table 3. In total, we identify 266 variants that contribute to AD variation substantially, which comes from 198 groups in 89 loci. Similar to the literature, multiple variants are identified in the APOE/APOC region with the strongest association to AD (chromosome 19, positions 44909011 $\sim$ 45912650). In addition, we also manage to identify variants rs6733839, rs744373 and rs2118506 close to gene “BIN1”, variant rs4844610 close to gene “CR1”, variants rs536841, rs10792832 and rs3844143 close to gene “PICALM”, variants rs4277405, rs4309 and rs3730025 close to gene “ACE”, all of which are also reported in He2022 and Bellenguez2022. Compared with the model-X knockoff filter, the FVG filter only misses 2 loci identified by the model-X knockoff filter but discover 53 more loci. The advantage of our FVG filter over KeLP is more dominating. No matter whether the intermediate layer is included in inference of KeLP, all KeLP-identified loci are discovered by our FVG filter while our FVG filter also returns at least another 44 new loci. This suggests that our method would not suffer from the same power loss issue of the model-X knockoff filter and KeLP. Compared to the group knockoff filter and KnockoffZoom [^4] KnockoffZoom’s result at the group-versus-group level is the same as the group knockoff filter, no matter whether the intermediate layer is included in inference. which identify 198 AD-associated variant groups from 91 loci under the same target FDR level $0.10$, our FVG filter does not suffer power loss. As shown in Figure 6 (a), among all variant groups identified by either the group knockoff filter (and KnockoffZoom) or our FVG filter [^5] Here, we say a variant group is identified by our FVG filter if it contains at least one variant identified by our FVG filter., 192 groups are identified by both methods while both our FVG filter and the group knockoff filter identify 6 groups not in common. This suggests the consistency of inference results between our FVG filter and the group knockoff filter (and KnockoffZoom). However, our FVG filter manages to detect signals at variant rs7225151 close to gene “LOC100130950”, while none of variant groups in the “LOC100130950” locus are identified by the group knockoff filter (and KnockoffZoom).
(a) Venn diagram of variant groups identified by group knockoff filter and our FVG filter.
(b) Manhattan plot of feature importance statistics of different genetic variant in a high-LD region between positions 31571218 $\sim$ 32682663, chromosome 6, under the FVG filter.
(c) Comparison of catching set size among different approaches.
(d) Comparison of catching set purity among different approaches.
Figure 6: Comparison of our FVG filter and other methods in analyzing the EADB–UKBB dataset.
Without power loss, our FVG filter returns informative selection sets via simultaneous fine-mapping. This can be seen in Figure 6 (c)-(d) where empirical distributions of catching set sizes and purity are displayed. Specifically, all but one catching sets of our FVG filter are of size no larger than $8$ while most of catching sets are of size 1 and purity greater than $90\%$. In other words, most of catching sets only contain the important variants or strongly correlated proxy variants. This results in the smallest average size (1.343) and highest average purity (0.951) of catching sets. In contrast, 32 catching sets obtained by the group knockoff filter contain more than $8$ variants with strong impurity and are far from being informative. Specifically, for the five largest catching sets obtained by the group knockoff filter shown in Table 4, our FVG filter manages to distinguish one or two variants from more than 50 proxy variants in each. For comparison, KnockoffZoom (2 layers) only manages to narrow the signal of group “6:31571218:32682663:64” down to the variant rs532965 while the rest four groups are included as a whole in the resolution-adaptive discoveries. Adding the intermediate layer only refines the catching set of group “17:43056905:45876021:36” (size: 314) to a catching set of subgroup with size 281 and the catching set of group “6:31571218:32682663:83” (size: 57) to a catching set of subgroup with size 19. Thus, the average size and purity of catching sets provided by KnockoffZoom is still suboptimal. In addition, catching sets provided by our FVG filter also dominate the KeLP in size. Specifically, among all catching sets provided by our FVG filter, more than 95% are if size 1 or 2 (less than 85% for KeLP (2 layers) and KeLP (3 layers)). This suggests the advantages of the proposed FVG filter in refining AD-associated genetic discoveries over KnockoffZoom and KeLP. Comprehensive comparisons of our method with all existing methods are summarized in Table 5.
@@
Identified Chromosome Position Catching sets obtained by Closest cS2Gene Annotation
@@
rs9271162 6 32609938 6:31571218:32682663:83 (57) HLA-DRB1 HLA-DRB1 EpiMap
rs9273349 6 32658092 HLA-DQB1
rs199451 17 46724418 17:43056905:45876021:36 (314) NSF KANSL1-AS1 GTeX
Table 5: Summary of results by applying our FVG filter to the EADB-UKBB dataset in comparison with existing methods, where bold values indicate the best performance under different evaluation metrics.
@@
Marginal association test 54 17.093 17.093 0.489
Model-X knockoff filter 38 4.263 1.000 1.000
Group knockoff filter 91 19.923 9.157 0.651
KnockoffZoom (2 layers) 91 15.152 5.621 0.851
KnockoffZoom (3 layers) 91 14.681 5.409 0.858
KeLP (2 layers) 45 10.422 2.535 0.918
KeLP (3 layers) 38 10.605 2.488 0.932
FVG filter 89 2.989 1.343 0.951
: Size and purity of catching sets provided by model-X knockoff filter are trivially 1 and thus not included
   in comparison.
In addition to statistical performance, the biological informativeness of discoveries made by our FVG filter is also our concern. We apply a SNP-to-gene linking strategy (Gazal2022) to investigate whether the identified genetic variants are functionally enriched or not. Here, genetic variants in the same group can functionally annotate different genes or have different annotation types. There are in total 7 types of annotations as follows.
  • Exon: The genetic variant is located less than 20 base pairs from an exon of the annotated gene.
  • Promoter: The genetic variant lies in promoter regions of the annotated gene.
  • GTEx: The genetic variant lies in an expression quantitative trait locus of the annotated gene across 54 cell types, according to the GTEx data.
  • eQTLGen: The genetic variant lies in an expression quantitative trait locus of the annotated gene in blood, according to the eQTLGen data.
  • Epimap: The genetic variant lies in enhancer regions of the annotated gene across 833 cell types under the EpiMap model.
  • ABC: The genetic variant lies in Hi-C linked enhancer regions of the annotated gene across 167 cell types under the Activity-by-Contact model.
  • Cicero: The genetic variant has co-accessibility links to the annotated gene’s promoter across 61,806 blood/basal cells.
Among 266 genetic variants identified by our FVG filter in Table 3, 180 (67.67%) can be mapped with function evidence. This proportion of functional annotation is significantly higher than the average percentage of the background genome (30.52%) and variants in catching sets of the group knockoff filter (56.54%) as shown in Table 6. Such an advantage is more significant in annotation types “Exon”, “Promoter” and “ABC”. As a result, compared to the group-level inference, discoveries made by the FVG filter is more biological informative for further drug development or therapy designs.
Table 6: Proportions of different annotations for variants in catching sets of group knockoff filter and the FVG filter.
Annotation Background Genome Group knockoff filter FVG filter
Exon 6.74% 11.64% 19.55%
Promoter 2.68% 7.67% 12.03%
GTeX 8.33% 22.78% 24.44%
eQTLGen 1.57% 6.56% 6.02%
Epimap 10.10% 12.24% 12.41%
ABC 8.87% 9.76% 14.66%
Cicero 0.98% 3.09% 4.51%
Total 30.52% 56.54% 67.67%